# Quantitative estimates: How well does the discrete Fourier transform approximate the Fourier transform on $\mathbb{R}$?

### v1

@@ 3.3 Polynomial weights
In this section we consider the polynomial weights $v_{\alpha}(x)=(1+|x|)^{\alpha}$. To bound $\Phi_{v_{\alpha}}(p)$ in (16), we make use of the Hurwitz zeta function $\zeta(s,t)=\sum_{m=0}^{\infty}(m+t)^{-s}$.
@@
$\Phi_{v_{\alpha}}(p)\leq p^{-\alpha}\sqrt{2\zeta(2\alpha,\tfrac{1}{2})}\leq p^{-\alpha}2^{\alpha+1/2}\Big{(}1+\frac{1}{4\alpha-2}\Big{)}^{1/2}\lesssim p^{-\alpha}\,.$
Proof 3.8.
@@
$\leq p^{-2\alpha}2\sum_{m=0}^{\infty}\left(m+\tfrac{1}{2}\right)^{-2\alpha}=p^{-2\alpha}2\zeta(2\alpha,\tfrac{1}{2})\,.$
The integral test for convergence of series provides, for $\alpha>1/2$,
@@
Corollary 3.9 (polynomial weights).
@@
    $E^{[n]}_{h}(f)\lesssim h^{\beta}\big{(}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\big{)}\,.$
    Expressed with the number of samples, the error is $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
Proof 3.10.
(i) The assumption $h\leq 1\leq p$ implies that $\sqrt{1+h}\leq\sqrt{2}$ and $\sqrt{1+p^{-1}}\leq\sqrt{2}$. By definition $v_{\alpha}(1)=2^{\alpha}$. Using Lemma 3.7, the constant $v_{\alpha}(1)\Phi_{v_{\alpha}}(p)\sqrt{1+h}$ in Theorem 3.5 is then bounded by $2^{2\alpha+1}(1+1/(4\alpha-2))^{1/2}$. Likewise for the constant involving $w=v_{\beta}$. The error estimate now follows from Theorem 3.5.
@@
Corollary 3.9 answers the questions (Q1) and (Q2) raised in the introduction whenever $f$ decays polynomially in time and in frequency.
@@
In the literature, e.g. [16], one can find the choice $h\asymp\frac{1}{\sqrt{n}}$. It is optimal only for $\alpha=\beta$. If $\alpha\neq\beta$, then $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to better bounds. Our analysis reveals that the spacing should depend on the decay of $f$ and $\hat{f}$, and Corollary 3.9 explains how to choose the optimal parameters.
In our experience, amalgam spaces are the natural theoretical framework when dealing with sampling and periodization. Yet one may wish to have conditions that are easier to check in practice. One such condition is a pure decay condition $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ as in Theorem 1.1.
Proof 3.12 (Proof of Theorem 1.1).
Assume that $|f(x)|\lesssim(1+|x|)^{-a}$ and $\alpha<a-\frac{1}{2}$. We show that $f\in W(\mathcal{C},\ell^{2}_{v_{\alpha}})$. Hölder’s inequality yields
$\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{2}(1+|l|)^{2\alpha}\lesssim\sup_{x\in\mathbb{R}}|f(x)|^{2}(1+|x|)^{2a}\sum_{l\in\mathbb{Z}}(1+|l|)^{-2(a-\alpha)}\,.$ (21)
@@ 3.4 Sub-exponential weights
Next, we consider sub-exponential weights of the form $v_{r,\alpha}(x)=e^{r|x|^{\alpha}}$ as defined in Section 3.2. We first find a bound for $\Phi_{v_{r,\alpha}}(p)$ in (16). We directly derive
$|\Phi_{v_{r,\alpha}}(p)|^{2}=2\sum_{m=0}^{\infty}e^{-2r\left(pm+\frac{p}{2}\right)^{\alpha}}=e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,.$ (22)
@@
$|\Phi_{v_{r,\alpha}}(p)|^{2}\leq e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{1}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,,$ (23)
@@
In particular, the factor $\big{(}2+\frac{2^{\alpha}}{2\alpha^{2}rp^{\alpha}}\big{)}^{1/2}$ tends to $\sqrt{2}$ when $p\rightarrow\infty$.
Proof 3.14.
@@
A direct check reveals that $-\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},x^{\alpha}u)$ is a primitive of $x\mapsto e^{-ux^{\alpha}}$ and therefore $\int_{1}^{\infty}e^{-ux^{\alpha}}\,\mathrm{d}x=-\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},u)$. Hence, we derive
$\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}\leq 1+\frac{1}{2\alpha}\frac{e^{u}}{u^{1/\alpha}}\Gamma(\tfrac{1}{\alpha},u)\,.$ (24)
@@
$\Gamma(\tfrac{1}{\alpha},u)\leq\frac{1}{\alpha}u^{\frac{1}{\alpha}-1}e^{-u}\,,$ (25)
see [18, Proposition 2.7 and Page 1275]. For completeness we reproduce the elementary argument:
Set $a=\alpha^{-1}$. The function $h(t)=(a-1)\ln t-t$ is strictly concave, so it is majorized by its tangent $h_{u}(t)=h(u)+h^{\prime}(u)(u-x)$. Therefore
[…]
@@
We substitute (25) into (24), use the assumption $u\leq 1/\alpha$, and obtain
@@
Proof 3.17 (Proof of Theorem 1.2).
Assume that $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $r^{\prime}=r-\epsilon<r$. Then $f\in W(\mathcal{C},\ell^{1}_{r-\epsilon,\alpha})$ for every $\epsilon>0$ with the same argument as in (21). Therefore the theorem is a direct consequence of Corollary 3.15 and this embedding.
@@ 3.5 Mixed weights
Assume that $f\in W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{\beta}})$. We feed the estimates of Lemma 3.7 and 3.13 into Theorem 3.5 and obtain
$E^{[n]}_{h}(f)\leq e^{r}\big{(}4+\tfrac{2}{\alpha}\big{)}^{1/2}e^{-r(\frac{p}{2})^{\alpha}}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+2^{2\beta+1}\big{(}1+\tfrac{1}{4\beta-2}\big{)}^{\frac{1}{2}}h^{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}$ (26)
@@
As for direct decay conditions, if $f$ and $\hat{f}$ are continuous and satisfy $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$ (27)
@@
For the exponential case $\alpha=1$, a related bound for the sup-norm of $\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}$ is derived by Briggs and Henson [4, Theorem 6.6]. However, they impose extremely restrictive regularity requirements. For a comparable bound the periodization $\mathcal{P}_{p}(f{\bf 1}_{[-\frac{p}{2},\frac{p}{2})]})$ must be $b-2$-times continuously differentiable on $\mathbb{R}$. This holds only for those interval lengths $p$, where the derivatives at ±$\pm\frac{p}{2}$ coincide.
@@ 4 Interpolation
So far, we have approximated the Fourier samples $\hat{f}(\tfrac{k}{p})$ by the discrete Fourier transform $h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}kj/n}=\sqrt{p}\,(\mathcal{F}f_{h,n})(k)$, for $k\in[n]$. In this section we change the point of view: we now interpolate the vector $\mathcal{F}f_{h,n}$ and study how the interpolating function approximates $\hat{f}$ on whole real line $\mathbb{R}$. To do so, we use three exemplary interpolation schemes that are often used in approximation theory.
The standard procedure starts with a cardinal interpolating function: this is a nice function $\phi$ on $\mathbb{R}$ that satisfies the interpolation condition $\phi(k)=\delta_{k}$. The approximation of $\hat{f}$ on $\mathbb{R}$ from the discrete Fourier transform is then given by
$\Phi_{h,n}(\xi)=\sqrt{p}\,\sum_{k\in[n]}\mathcal{F}f_{h,n}(k)\phi(p\xi-k)=h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\phi(p\xi-k)\,.$ (28)
This expression is an approximation of $\hat{f}$ on $\mathbb{R}$ that satisfies $\Phi_{h,n}(\tfrac{k}{n})=\sqrt{p}\,\mathcal{F}f_{h,n}(k)$, and we seek to estimate the global error $\|\hat{f}-\Phi_{h,n}\|_{2}$.
As the cardinal interpolating function $\phi$ we will use the first two B-splines $B_{1}={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$ and $B_{2}(x)=(1-|x|){\bf 1}_{[-1,1]}(x)$ and the cardinal sine function $\operatorname*{sinc}(x)=\tfrac{\sin\pi x}{\pi x}$. Clearly, these functions satisfy $\phi(k)=\delta_{k}$. To keep the notion compact, we set $B_{0}:=\operatorname*{sinc}$. For $\phi=B_{1}$, then right-hand side of (28) is a step function, for $\phi=B_{2}$ we obtain a continuous, piecewise linear function (this is how many software would plot the approximating vector $\mathcal{F}f_{h,n}$), and for $\phi=\operatorname*{sinc}$ we obtain a smooth function whose Fourier transform has support in $[-1/2,1/2]$, which is commonly referred to as a bandlimited function.
4.1 Convergence rates
We first derive an error bound for the bandlimited approximation (28) with $\Phi=\operatorname*{sinc}$ for general weights. This is then specified to the polynomial and the sub-exponential weights, resulting in explicit error rates. The approximation rates for the $B$-splines (28) with $i=1,2$ are restricted to $\alpha\leq 1$ and $\alpha\leq 2$, respectively.
@@ 4.1 Convergence rates
Assume that $0<h\leq 1\leq p$ and that the weights $v,w$ satisfy (3.2). If $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$, then
$\Big{\|}\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (29)
If $\operatorname*{sinc}$ is replaced with $B_{i}$, for $i=1,2$, then the estimate (29) bounds holds for the polynomial weights $v=v_{\alpha}$ subject to the restriction $1/2<\alpha\leq i$.
We note that the error is of the same form as in Theorem 3.5. In particular, for polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$ with $\alpha,\beta>\frac{1}{2}$ we obtain $\Phi_{v}(p)\lesssim p^{-\alpha}$ and $\Phi_{w}(h^{-1})\lesssim h^{\beta}$ (see Lemma 3.7). For sub-exponential weights $v=v_{r,\alpha}$ and $w=v_{s,\beta}$ with $0<\alpha,\beta\leq 1$ and $r,s>0$ we obtain $\Phi_{v}(p)\lesssim e^{-r(\frac{p}{2})^{\alpha}}$ and $\Phi_{w}(h^{-1})\lesssim e^{-s(\frac{1}{2h})^{\beta}}$ (see Lemma 3.13).
Proof 4.2.
@@
$\hat{f}\,\overset{(s1)}{\approx}\,\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)$ $\,\overset{(s2)}{\approx}\,\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)$
$\,\overset{(s3)}{\approx}\,h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\cdot-k)\,.$ (30)
We derive bounds for the steps (s2) and (s3) with respect to general weights simultaneously for $i=0,1,2$.
(s3) For $i=0,1$, the system $\{\sqrt{p}B_{i}(p\cdot-k)\}_{k\in\mathbb{Z}}$ is orthonormal in $L^{2}(\mathbb{R})$ and it is still a Riesz sequence for $i=2$ with Riesz bounds independent of $p$ (this can be verified by direct computation). Therefore, we conclude that
$\Big{\|}\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\cdot-k)\Big{\|}_{L^{2}}$
$=\|\sum_{k\in[n]}\frac{1}{\sqrt{p}}\Big{(}\hat{f}(\frac{k}{p})-\mathcal{F}f_{h,n}(k)\Big{)}\,\sqrt{p}\,B_{i}(p\cdot-k)\|_{L}^{2}\asymp\big{\|}\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\big{\|}\ =E^{[n]}_{h}(f)\,,$
@@
$\Big{\|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}^{2}_{L^{2}}\asymp p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}\,.$
@@
The sum on the right-hand side is bounded as in estimates at the end of the proof of Lemma 3.3. We deduce

### v2

@@ 4.5 Polynomial weights
In this section we consider the polynomial weights $v_{\alpha}(x)=(1+|x|)^{\alpha}$ and derive bounds on $\Phi_{v_{\alpha}}(p)$ in (24).
@@
$\Phi_{v_{\alpha}}(p)\leq p^{-\alpha}2^{\alpha+1/2}\Big(1+\frac{1}{4\alpha-2}\Big)^{1/2}\lesssim p^{-\alpha}\,.$
Proof 4.10.
@@
$\leq p^{-2\alpha}2\sum_{m=0}^{\infty}\left(m+\tfrac{1}{2}\right)^{-2\alpha}\,.$
This sum has the shape of the Hurwitz zeta function $\zeta(s,t)=\sum_{m=0}^{\infty}(m+t)^{-s}$ at $t=1/2$. The integral test for convergence of series provides, for $\alpha>1/2$,
@@
Theorem 4.11 (polynomial weights).
@@
    $E^{[n]}_{h}(f)\lesssim h^{\beta}\big(\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\big)\,$,
    where the constant in $\lesssim$ is independent of $h,n,p,f$, but may depend on $\alpha,\beta$. Expressed with the number of samples, the error is
    $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$
    with a constant independent of $h,n,p$.
Part (ii) can be made more explicit by collecting all constants, leading to the ideal choice of $h$ in relation to properties of $f$.
Corollary 4.12 (explicit constants).
Under the assumptions of Theorem 4.11, we make the choice
$h=\left(\frac{c_{\alpha}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}}{c_{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}}\right)^{\frac{1}{\alpha+\beta}}n^{-\frac{\alpha}{\alpha+\beta}}\,$,
where $c_{s}=2^{2s+1}\left(1+\tfrac{1}{4s-2}\right)^{\frac{1}{2}}$. Then the resulting error bound becomes
$E^{[n]}_{h}(f)\leq 2\left(c_{\alpha}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}\right)^{\frac{\beta}{\alpha+\beta}}\left(c_{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\right)^{\frac{\alpha}{\alpha+\beta}}n^{-\frac{\alpha\beta}{\alpha+\beta}}\,.$
[…]
Proof 4.13.
(i) The assumption $h\leq 1\leq p$ implies that $\sqrt{1+h}\leq\sqrt{2}$ and $\sqrt{1+p^{-1}}\leq\sqrt{2}$. We note $v_{\alpha}(1)=2^{\alpha}$. Using Lemma 4.9, the constant $v_{\alpha}(1)\Phi_{v_{\alpha}}(p)\sqrt{1+h}$ in Theorem 4.7 is then bounded by $2^{2\alpha+1}(1+1/(4\alpha-2))^{1/2}$, likewise for the constant involving $w=v_{\beta}$. The error estimate now follows from Theorem 4.7.
@@
Theorem 4.11 and Corollary 4.12 answer the questions (Q1) and (Q2) raised in the introduction whenever $f$ decays polynomially in time and in frequency.
@@
In the literature, e.g. [4, 20], one can find the choice $h\asymp\frac{1}{\sqrt{n}}$. It is optimal for $\alpha=\beta$. If $\alpha\neq\beta$, then $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to better bounds. That is, knowledge on the decay of $f$ and $\hat{f}$ allows us to refine the spacing, and Theorem 4.11 explains how to choose the optimal parameters.
In our experience, amalgam spaces are the natural theoretical framework when dealing with sampling and periodization. Yet one may wish to have conditions that are easier to check in practice. One such condition is a pure decay condition of the form $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ as in Theorem 2.1.
Proof 4.15 (Proof of Theorem 2.1).
Assume that $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ and $\alpha<a-\frac{1}{2}$. We show that $f\in W(\mathcal{C},\ell^{2}_{v_{\alpha}})$. Hölder’s inequality yields
(29) $\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{2}(1+|l|)^{2\alpha}\lesssim\sup_{x\in\mathbb{R}}|f(x)|^{2}(1+|x|)^{2a}\sum_{l\in\mathbb{Z}}(1+|l|)^{-2(a-\alpha)}\,.$
@@ 4.6 Sub-exponential weights
Next, we consider sub-exponential weights of the form $v_{r,\alpha}(x)=e^{r|x|^{\alpha}}$. We first find a bound for $\Phi_{v_{r,\alpha}}(p)$ in (24). We directly derive
(30) $|\Phi_{v_{r,\alpha}}(p)|^{2}=2\sum_{m=0}^{\infty}e^{-2r\left(pm+\frac{p}{2}\right)^{\alpha}}=e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,.$
@@
(31) $|\Phi_{v_{r,\alpha}}(p)|^{2}\leq e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{1}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,$,
@@
Note that the factor $\big(2+\frac{2^{\alpha}}{2\alpha^{2}rp^{\alpha}}\big)^{1/2}$ tends to $\sqrt{2}$ when $p\rightarrow\infty$.
Proof 4.17.
@@
A direct check reveals that $\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},x^{\alpha}u)$ is a primitive of $x\mapsto e^{-ux^{\alpha}}$ and therefore $\int_{1}^{\infty}e^{-ux^{\alpha}}\,\mathrm{d}x=\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},u)$. Hence, we derive
(32) $\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}\leq 1+\frac{1}{2\alpha}\frac{e^{u}}{u^{1/\alpha}}\Gamma(\tfrac{1}{\alpha},u)\,.$
@@
(33) $\Gamma(\tfrac{1}{\alpha},u)\leq\frac{1}{\alpha}u^{\frac{1}{\alpha}-1}e^{-u}\,$,
see [23, Proposition 2.7 and Page 1275]. For completeness we reproduce the elementary argument: set $a=\alpha^{-1}$. The function $h(t)=(a-1)\ln t-t$ is strictly concave, so it is majorized by its tangent $h_{u}(t)=h(u)+h^{\prime}(u)(t-u)$ at $u$. Therefore
$\Gamma(a,u)=\int_{u}^{\infty}e^{h(t)}\,dt<\int_{u}^{\infty}e^{h_{u}(t)}\,dt=\frac{u^{a-1}e^{-u}}{1-(a-1)/u}\leq a{u^{a-1}e^{-u}}\,$,
@@
We substitute (33) into (32), use the assumption $u\geq 1/\alpha$, and obtain
@@
Proof 4.20 (Proof of Theorem 2.3).
Assume that $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}<\infty$ and $r^{\prime}=r-\epsilon<r$. Then $f\in W(\mathcal{C},\ell^{2}_{r-\epsilon,\alpha})$ for every $\epsilon>0$ with the same argument as in (29). Therefore the theorem is a consequence of Corollary 4.18.
@@ 4.7 Mixed weights
Assume that $f\in W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{\beta}})$. We feed the estimates of Lemma 4.9 and 4.16 into Theorem 4.7 and obtain
(34) $E^{[n]}_{h}(f)\leq e^{r}\big(4+\tfrac{2}{\alpha}\big)^{1/2}e^{-r(\frac{p}{2})^{\alpha}}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+2^{2\beta+1}\big(1+\tfrac{1}{4\beta-2}\big)^{1/2}h^{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}$
@@
As for direct decay conditions, if $f$ and $\hat{f}$ are continuous and satisfy
$\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
(35) $E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$
@@
For the exponential case $\alpha=1$, a related bound for the sup-norm of $\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}$ is derived by Briggs and Henson [5, Theorem 6.6]. However, they impose restrictive boundary conditions. For a comparable bound the periodization $\mathcal{P}_{p}(f{\bf 1}_{[-\frac{p}{2},\frac{p}{2})]})$ must be b2b-2-times continuously differentiable on $\mathbb{R}$. These are already violated when $f(p/2)\neq f(-p/2)$, in which case (35) holds only with $\beta=1$ regardless of the global smoothness of $f$.
@@ 5 Interpolation
So far, we have approximated the Fourier samples $\hat{f}(\tfrac{k}{p})$ by the discrete Fourier transform $h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}kj/n}=\sqrt{p}\,(\mathcal{F}f_{h,n})(k)$, for $k\in[n]$. In this section we change the point of view: we now interpolate the vector $\sqrt{p}\,\mathcal{F}f_{h,n}$ and study how the interpolating function approximates $\hat{f}$ on whole real line $\mathbb{R}$. To do so, we use three exemplary interpolation schemes that are often used in approximation theory, namely piecewise constant and piecewise linear interpolation, and interpolation using sinc functions.
The standard procedure starts with a cardinal interpolating function: this is a nice function $\phi$ on $\mathbb{R}$ that satisfies the interpolation condition $\phi(k)=\delta_{k,0}$. The approximation of $\hat{f}$ on $\mathbb{R}$ from the discrete Fourier transform is then given by
(36) $\Psi_{h,n}(\xi)=\sqrt{p}\,\sum_{k\in[n]}\mathcal{F}f_{h,n}(k)\phi(p\xi-k)=h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\phi(p\xi-k)\,.$
This expression is an approximation of $\hat{f}$ on $\mathbb{R}$ that satisfies $\Psi_{h,n}(\tfrac{k}{p})=\sqrt{p}\,\mathcal{F}f_{h,n}(k)$, and we seek to estimate the global error $\|\hat{f}-\Psi_{h,n}\|_{L^{2}}$.
As the cardinal interpolating function $\phi$ we will use the first two B-splines $B_{1}={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$ and $B_{2}(x)=(1-|x|){\bf 1}_{[-1,1]}(x)$ and the cardinal sine function $\operatorname*{sinc}(x)=\tfrac{\sin\pi x}{\pi x}$. Clearly, these functions satisfy $\phi(k)=\delta_{k,0}$. To keep the notion compact, we set $B_{0}:=\operatorname*{sinc}$. For $\phi=B_{1}$, the right-hand side of (36) is a step function, for $\phi=B_{2}$ we obtain a continuous, piecewise linear function (this is how many software programs would plot the approximating vector $\mathcal{F}f_{h,n}$), and for $\phi=\operatorname*{sinc}$ we obtain a smooth function whose Fourier transform has support in [1/2,1/2][-1/2,1/2], which is commonly referred to as a bandlimited function in time-frequency analysis.
We first derive an error bound for the bandlimited approximation (36) with $\phi=\operatorname*{sinc}$ for general weights. This is then specialized to the polynomial and the sub-exponential weights, resulting in explicit error rates. The approximation rates for the $B$-splines (36) with $i=1,2$ are restricted to decay rates $\alpha\leq 1$ and $\alpha\leq 2$, respectively.
@@
Assume that $0<h\leq 1\leq p$ and that the weights $v,w$ satisfy (22). If $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$, then
(37) $\hskip-4.30554pt\Big\|\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
If $\operatorname*{sinc}$ is replaced with $B_{i}$, for $i=1,2$, then the estimate (37) still holds for the polynomial weight $v=v_{\alpha}$ subject to the restriction $1/2<\alpha\leq i$.
We note that the error is of the same form as in Theorem 4.7. In particular, (i) for the polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$ with $\alpha,\beta>\frac{1}{2}$ we obtain $\Phi_{v}(p)\lesssim p^{-\alpha}$ and $\Phi_{w}(h^{-1})\lesssim h^{\beta}$ (see Lemma 4.9), and the choice $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to
$\Big\|\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}\left(\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\right)\,.$
(ii) For sub-exponential weights $v=v_{r,\alpha}$ and $w=v_{s,\beta}$ with $0<\alpha,\beta\leq 1$ and $r,s>0$ we obtain $\Phi_{v}(p)\lesssim e^{-r(\frac{p}{2})^{\alpha}}$ and $\Phi_{w}(h^{-1})\lesssim e^{-s(\frac{1}{2h})^{\beta}}$ (see Lemma 4.16).
(iii) For exponential weights $v(x)=e^{r|x|},w(x)=e^{s|x|}$ and the choice $p=\sqrt{n},h=1/\sqrt{n}$ we obtain the root-exponential convergence
$\Big\|\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim e^{-d\sqrt{n}}\,$,
as in  [28].
Proof 5.2 (Proof of Theorem 5.1).
@@
$\hat{f}(\xi)\,\overset{(s1)}{\approx}\,\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$ $\,\overset{(s2)}{\approx}\,\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$
(38) $\,\overset{(s3)}{\approx}\,h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\,.$
Here (s1) is the interpolation error, (s2) the truncation error, and (s3) the discretization error. We derive bounds for the error in the steps (s2) and (s3) with respect to general weights simultaneously for $i=0,1,2$.
(s3) For $i=0,1$, the system $\{\sqrt{p}B_{i}(p\xi-k)\}_{k\in\mathbb{Z}}$ is orthonormal in $L^{2}(\mathbb{R})$ and it is still a Riesz sequence for $i=2$ with Riesz bounds independent of $p$ (this can be verified by direct computation). Therefore, we conclude that
$\Big\|\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\Big\|_{L^{2}}$
$=\Big\|\sum_{k\in[n]}\Big(\tfrac{1}{\sqrt{p}}\hat{f}(\tfrac{k}{p})-\mathcal{F}f_{h,n}(k)\Big)\,\sqrt{p}\,B_{i}(p\xi-k)\Big\|_{L^{2}}\asymp\big\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\big\|\ =E^{[n]}_{h}(f)\,$,
@@
$\Big\|\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big\|^{2}_{L^{2}}\asymp p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}\,.$
@@
The sum on the right-hand side is bounded as in the estimates at the end of the proof of Lemma 4.5. We deduce